OpenStudy (anonymous):

Help please?? If you deposit $750 in an account that pays 8% annual interest compounded continuously, what will the balance be after five years? $29,803.91 $1,832.06 $1,277.16 $1,118.87

OpenStudy (mathmale):

For continuous compounding, A=Pe^(rt), where P=initial dollar amount; r=interest rate expressed as a decimal fraction; and t=time (in years). Is this sufficient info to enable you to calculate A (the amount the account owner will have in the account after 5 years?

OpenStudy (anonymous):

The equation that you would want is this, y=750(1.08x) The 750 represents your original investment, the 1.08 would represent that for each year (x) you get paid 8% interest. That would be the 1.08 part.

OpenStudy (mathmale):

Mason: I stick with my original statement: for continuous compounding, A=Pe^(rt). For annual compounding (payment of interest once per year), A=P(1.08)^t when the interest rate happens to be 8%; t represents the number of years.

OpenStudy (anonymous):

Just putting it in words that might be easier to understand for them. Im not saying your wrong, just another way to say it.

OpenStudy (anonymous):

@DaBaller with the equation I am getting 4,050 that is not one of my answers though.. I'm confused -_-

OpenStudy (anonymous):

ya, actually my bad, one sec i messed up on my typing... ill get you the real equation

OpenStudy (anonymous):

here it is... y=750(1.08)^5 everything represents the same as before, except the ^5 is the # of years. but after looking at it closer, i realized that this would not match up with your equation either, i believe your options are written wrong, or i am totally drunk...

OpenStudy (mathmale):

The ball's in your court, DaBaller. Pretty please look up "continuous compounding." Then decide whether or not I'm still "wrong." Continuous compounding implies that interest is being paid all the time, without interruptions; annual compounding implies that interest is paid once per year and added to the principal, and that interest is paid on this combined principal and interest at the end of the 2nd year, and so on, and so on. These two situations are NOT the same.

OpenStudy (anonymous):

never said you were wrong, and like i said, im probably drunk

OpenStudy (anonymous):

guess ill just go back to balling :)

OpenStudy (mathmale):

With continuous compounding, the Saver in Emm's question will accumulate $1,118.87 in five years. With annual compounding, somebody else will accumulate $750(1.08)^5, or $1,102.00, in five years. Glad you apparently have seen the light.